a tennis ball is hit vertically upwards with a velocity of 20m/s. calculate it's retardation as it moves up and the maximum height of the ball
1
Expert's answer
2020-03-17T09:51:41-0400

Retardation:


"a=-g=-9.8\frac{m}{s^2}"

The maximum height of the ball:


"h=\frac{v^2}{2g}=\frac{20^2}{2(9.8)}=20\ m"